OpenStudy (anonymous):

Need help finding the parallel lines to y=-3x-6;(-1,5)

OpenStudy (dan815):

now u just rearrange that in y=mx+b form to make it look nice

OpenStudy (dan815):

sorry a mistake that shud be

OpenStudy (dan815):

you see -3 must be a slope so it is parallel and you can write a line in this form so it goes through this point given like so y-5=-3(x-(-1)) y-5=-3(x+1) y-5=-3x-3 y=-3x+2

OpenStudy (dan815):

to make it more clear the general form is

OpenStudy (dan815):

if it goes through the time (a,b) and has to be parallel to line with slope m, then y-b=m(x-a)

OpenStudy (anonymous):

ok thanks

OpenStudy (dan815):

there are other ways to do it too

OpenStudy (dan815):

like say know the line intetrcept form y=mx+b

OpenStudy (dan815):

u know it has slope -3 so y=-3x+b now sub in that point we are given to find what b must be 5=-3(-1)+b b=5-3=2 so y=-3x+2

OpenStudy (dan815):

this is called the slope intercept form. m= slope and b= y-intercept

OpenStudy (anonymous):

ok not I have to find the perpendicular lines to y=1/3x+1;(4,2)

OpenStudy (dan815):

for a perpendicular line, you have to take the negative reciprocal of the slope

OpenStudy (dan815):

and repeat the same steps to find the b, for that given point

OpenStudy (dan815):